OpenStudy (anonymous):

A group holds a raffle to raise money. They sell 100 raffle tickets for $5 each. There is one grand prize worth $100, three second place prizes worth $25 each and 5 3rd place prizes worth $10 each. Find the expected value associated with the purchase of one ticket.

OpenStudy (anonymous):

anyone? :/

OpenStudy (king):

u want to find the profit per ticket earned rite?

OpenStudy (anonymous):

@UnkleRhaukus can you please help me?

OpenStudy (unklerhaukus):

you are looking for the expected return value of a ticket

OpenStudy (anonymous):

can you help me with the question though?

OpenStudy (unklerhaukus):

total ticket purchases are $5x100=$ total returns are 1x$100+3x$25+5x$10=$

OpenStudy (king):

thatz 225$

OpenStudy (unklerhaukus):

the expected value will be\[ $5\times\frac{$225}{$500} \]

OpenStudy (king):

2.25$?

OpenStudy (king):

thats the profit per ticket....thats what i said!

OpenStudy (unklerhaukus):

no, $2.25 is what each ticket is worth, the profit per ticket will be $5-$2.25=$2.75

OpenStudy (unklerhaukus):

i am not sure if the question is asking for the excepted value for the ticket or for the expected value of a sale of a ticket

OpenStudy (anonymous):

I think you're right $2.75